### v3

# Gauge-Invariant Entire-Function Regulators and UV Finiteness in Non-Local Quantum Field Theory

Abstract
We clarify the status of gauge-invariant entire-function regulators in NonLocal Quantum Field Theory. The regulator is implemented as an entire function of the covariant Laplace–Beltrami operator. Working in the background-field formalism and expanding around flat, trivial backgrounds, we show that plane waves diagonalize the d’Alembertian, so that the entire function reduces to a multiplicative form factor in Minkowski momentum space. After Wick rotation, to the Euclidean axis, producing exponential ultraviolet damping in loop integrals without introducing additional poles or branch cuts. Our analysis provides a concise, gauge-covariant justification for the use of entire-function regulators in nonlocal quantum field theory.
I Entire-function regulators and spectral calculus
We begin on a general Lorentzian manifold $(M,g_{\mu\nu})$, where $M$ is a smooth four-dimensional spacetime manifold and $g_{\mu\nu}$ is a Lorentzian metric with signature (,+,+,+)(-,+,+,+). We couple matter fields to a gauge field taking values in the Lie algebra $\mathfrak{g}$ of a gauge group $G$, and introduce a principal $G$-bundle $P\to M$. Matter fields transform in some representation of $G$.
The gauge- and diffeomorphism-covariant derivative $D_{\mu}$ acting on a field $\Phi$ is defined by [^1] There are past publications on nonlocal QFT. We provide a list of references for these publications in [9, 10, 7, 8, 4, 5, 6, 2, 3, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 26, 24, 25]:
$D_{\mu}\Phi\;\equiv\;\nabla^{\text{LC}}_{\mu}\Phi+A_{\mu}\Phi,$ (1)
where $\nabla^{\text{LC}}_{\mu}$ is the Levi–Civita covariant derivative associated with the metric $g_{\mu\nu}$, and $A_{\mu}$ is the gauge connection one-form in the appropriate representation of $\mathfrak{g}$. The covariant Laplace–Beltrami operator, also referred to as the Bochner d’Alembertian in this context, is then given by:
$\Box\;\equiv\;-g^{\mu\nu}D_{\mu}D_{\nu},$ (2)
where $g^{\mu\nu}$ is the inverse of the metric $g_{\mu\nu}$. When acting on a scalar field $\phi$, this reduces to:
$\Box\phi=-g^{\mu\nu}\nabla^{\text{LC}}_{\mu}\nabla^{\text{LC}}_{\nu}\phi.$ (3)
The operator $D_{\mu}\to\nabla^{\text{LC}}_{\mu}$ assumes the scalar field is a gauge singlet, so the gauge field $A_{\mu}$ acts trivially and drops out. If $\phi$ were charged, the fully gauge-covariant box operator would instead be:
$\Box\phi$ $=-g^{\mu\nu}D_{\mu}D_{\nu}\phi$ (4)
$=-g^{\mu\nu}(\nabla^{\text{LC}}_{\mu}+A_{\mu})(\nabla^{\text{LC}}_{\nu}+A_{\nu})\phi.$ (5)
Now, we let $F(z)$ be an entire function of a complex variable $z$. For definiteness, we will often take:
$F(z)=e^{z},$ (6)
the arguments below apply to any entire function $F$ with no zeros in the finite complex plane.
The operator $F(\Box/M_{*}^{2})$ is then defined by the convergent power series:
$F\!\left(\frac{\Box}{M_{*}^{2}}\right)\;\equiv\;\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{\Box}{M_{*}^{2}}\right)^{n},$ (7)
where $M_{*}$ is a constant mass scale that sets the onset of nonlocality. To understand the action of $F(\Box/M_{*}^{2})$, it is useful to recall the spectral calculus for self-adjoint operators. On a globally hyperbolic spacetime one can define a suitable Hilbert space of fields, and at least in the Euclidean or Riemannian case the analogue of $\Box$ is an essentially self-adjoint operator. We then have a spectral decomposition:
$\Box\psi_{\lambda}=\lambda\psi_{\lambda},$ (8)
where $\psi_{\lambda}$ is an eigenfunction with eigenvalue $\lambda$. Acting on such an eigenfunction, the operator $F(\Box/M_{*}^{2})$ reduces to multiplication by the scalar $F(\lambda/M_{*}^{2})$:
$F\!\left(\frac{\Box}{M_{*}^{2}}\right)\psi_{\lambda}=F\!\left(\frac{\lambda}{M_{*}^{2}}\right)\psi_{\lambda}.$ (9)
In curved spacetime with nontrivial gauge backgrounds, the spectrum of $\Box$ is complicated. However, to derive perturbative Feynman rules it suffices to expand around the perturbative vacuum, where the background is flat and the gauge connection vanishes. In that case the eigenfunctions become plane waves, and the spectrum becomes the familiar Minkowski momentum spectrum. This is the regime in which we will make the connection to Euclidean momentum space explicit.
II Flat-space limit and plane-wave eigenfunctions
To specialize to the perturbative vacuum, we take:
$g_{\mu\nu}\to\eta_{\mu\nu}=\text{diag}(-1,1,1,1),\qquad A_{\mu}\to 0,$ (10)
where $\eta_{\mu\nu}$ is the Minkowski metric and $A_{\mu}$ is the gauge connection. In this background $D_{\mu}$ reduces to the ordinary derivative:
$D_{\mu}\to\partial_{\mu}\equiv\frac{\partial}{\partial x^{\mu}},$ (11)
and the covariant box (2) reduces to the familiar Minkowski d’Alembertian:
$\Box\to\Box_{M}\equiv-\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}=-\left(-\partial_{0}^{2}+\partial_{i}\partial_{i}\right)=\partial_{0}^{2}-\nabla^{2}.$ (12)
where $\partial_{0}=\partial/\partial t$ is the derivative with respect to time $t=x^{0}$, $\partial_{i}=\partial/\partial x^{i}$ for $i=1,2,3$ are the spatial derivatives, and $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$ is the spatial Laplacian.
We consider a scalar field $\phi(x)$ on Minkowski spacetime $\mathbb{R}^{1,3}$. We define its Fourier transform by:
$\phi(x)=\int\frac{d^{4}p}{(2\pi)^{4}}\,e^{-ip\cdot x}\,\tilde{\phi}(p),$ (13)
where $p^{\mu}=(E,\bm{p})$ is the Minkowski four-momentum, with energy component $E=p^{0}$ and spatial momentum $\bm{p}=(p^{1},p^{2},p^{3})$, $p\cdot x\equiv p_{\mu}x^{\mu}=-Et+\bm{p}\cdot\bm{x}$ is the Minkowski inner product of $p^{\mu}$ and $x^{\mu}$, and $\tilde{\phi}(p)$ is the momentum-space field. Acting $\Box_{M}$ on a plane wave $e^{-ip\cdot x}$ gives:
$\partial_{0}e^{-ip\cdot x}$ $=\frac{\partial}{\partial t}e^{-i(-Et+\bm{p}\cdot\bm{x})}=iE\,e^{-ip\cdot x},$ (14)
$\partial_{i}e^{-ip\cdot x}$ $=\frac{\partial}{\partial x^{i}}e^{-i(-Et+p_{j}x^{j})}=-ip_{i}\,e^{-ip\cdot x},$ (15)
where the repeated spatial index $j$ is summed from $1$ to $3$. Differentiating once more, we find:
$\partial_{0}^{2}e^{-ip\cdot x}$ $=\partial_{0}(iEe^{-ip\cdot x})=iE(iE)e^{-ip\cdot x}=-E^{2}e^{-ip\cdot x},$ (16)
$\partial_{i}^{2}e^{-ip\cdot x}$ $=\partial_{i}(-ip_{i}e^{-ip\cdot x})=-ip_{i}(-ip_{i})e^{-ip\cdot x}=-p_{i}^{2}e^{-ip\cdot x}.$ (17)
Summing over spatial indices with $\delta^{ij}$ gives:
$\nabla^{2}e^{-ip\cdot x}=\delta^{ij}\partial_{i}\partial_{j}e^{-ip\cdot x}=-\bm{p}^{2}e^{-ip\cdot x},$ (18)
where $\bm{p}^{2}\equiv\delta^{ij}p_{i}p_{j}$. Now using the definition (12), we obtain:
$\Box_{M}e^{-ip\cdot x}$ $=(\partial_{0}^{2}-\nabla^{2})e^{-ip\cdot x}$
$=\left(-E^{2}-(-\bm{p}^{2})\right)e^{-ip\cdot x}$
$=(E^{2}-\bm{p}^{2})e^{-ip\cdot x}.$ (19)
It is standard to define the Lorentz-invariant norm of the momentum by:
$p^{2}\equiv\eta_{\mu\nu}p^{\mu}p^{\nu}=-E^{2}+\bm{p}^{2},$ (20)
so that:
$E^{2}-\bm{p}^{2}=-p^{2}.$ (21)
We then arrive at the eigenvalue equation:
$\Box_{M}e^{-ip\cdot x}=-p^{2}e^{-ip\cdot x}.$ (22)
By linearity, if we insert the Fourier expansion (13) into $\Box_{M}\phi(x)$, we find:
$\Box_{M}\phi(x)$ $=\int\frac{d^{4}p}{(2\pi)^{4}}\Box_{M}\left(e^{-ip\cdot x}\right)\tilde{\phi}(p)$
$=\int\frac{d^{4}p}{(2\pi)^{4}}\left(-p^{2}e^{-ip\cdot x}\right)\tilde{\phi}(p).$ (23)
So, in momentum space, $\Box_{M}$ acts simply as multiplication by $p^{2}$. Now we will consider the entire function of $\Box_{M}$ defined by (7) and (6):
$F\!\left(\frac{\Box_{M}}{M_{*}^{2}}\right)=\exp\!\left(\frac{\Box_{M}}{M_{*}^{2}}\right)=\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{\Box_{M}}{M_{*}^{2}}\right)^{n}.$ (24)
Acting on a plane wave, each power of $\Box_{M}$ contributes a factor $p^{2}$:
$\left(\frac{\Box_{M}}{M_{*}^{2}}\right)^{n}e^{-ip\cdot x}=\left(\frac{-p^{2}}{M_{*}^{2}}\right)^{n}e^{-ip\cdot x}.$ (25)
Summing over $n$, we obtain:
$F\!\left(\frac{\Box_{M}}{M_{*}^{2}}\right)e^{-ip\cdot x}$ $=\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{-p^{2}}{M_{*}^{2}}\right)^{n}e^{-ip\cdot x}$
$=\exp\!\left(-\frac{p^{2}}{M_{*}^{2}}\right)e^{-ip\cdot x}.$ (26)
This is precisely the statement specialized to the choice $F(z)=e^{z}$.
III From Minkowski to Euclidean momentum space
The representation (26) is the starting point for understanding the effect of the regulator in Feynman loop integrals. For a full picture, consider a scalar field $\phi$ of mass $m$ with a regulated propagator in Minkowski space. The ordinary Feynman propagator is:
$\Delta_{F}(p)=\frac{i}{p^{2}-m^{2}+i\epsilon},$ (27)
where $\epsilon>0$ is an infinitesimal parameter specifying the Feynman contour. If we dress the propagator with the entire-function regulator, we obtain:
$\Delta_{F}^{\text{reg}}(p)=\frac{i\,F(-p^{2}/M_{*}^{2})}{p^{2}-m^{2}+i\epsilon}.$ (28)
For $F(z)=e^{z}$, this becomes:
$\Delta_{F}^{\text{reg}}(p)=\frac{i\,e^{-p^{2}/M_{*}^{2}}}{p^{2}-m^{2}+i\epsilon}.$ (29)
A typical one-loop contribution, to a two-point function in $\phi^{4}$ theory would then involve integrals of the form:
$I=\int\frac{d^{4}p}{(2\pi)^{4}}\,\frac{i\,e^{-p^{2}/M_{*}^{2}}}{p^{2}-m^{2}+i\epsilon},$ (30)
where $p^{\mu}$ is the loop momentum integrated over Minkowski momentum space. To make the UV convergence properties manifest and to connect to the Euclidean damping factor, we perform a Wick rotation. We define the Euclidean four-momentum $p_{E}^{\mu}$ by:
$p^{0}=E\;\to\;ip_{E}^{0},\qquad\bm{p}\to\bm{p}_{E},$ (31)
where $\bm{p}_{E}$ is identified with the same numerical components as $\bm{p}$, but now interpreted as Euclidean spatial components. The Euclidean norm of $p_{E}^{\mu}=(p_{E}^{0},\bm{p}_{E})$ is:
$p_{E}^{2}\equiv\delta_{\mu\nu}p_{E}^{\mu}p_{E}^{\nu}=(p_{E}^{0})^{2}+\bm{p}_{E}^{2}.$ (32)
Under the Wick rotation (31), the Minkowski invariant (20) transforms as:
$p^{2}$ $=-E^{2}+\bm{p}^{2}$
$\to-(-p_{E}^{0})^{2}+\bm{p}_{E}^{2}$
$=(p_{E}^{0})^{2}+\bm{p}_{E}^{2}$
$=p_{E}^{2}.$ (33)
Thus $p^{2}\to p_{E}^{2}$. The integration measure transforms as:
$dp^{0}=i\,dp_{E}^{0},\qquad d^{4}p=dp^{0}\,d^{3}\bm{p}=i\,dp_{E}^{0}\,d^{3}\bm{p}_{E}=i\,d^{4}p_{E}.$ (34)
Inserting these relations into (30) and using the standard contour deformation argument, we obtain a Euclidean integral:
$I_{E}=\int\frac{d^{4}p_{E}}{(2\pi)^{4}}\,\frac{e^{-p_{E}^{2}/M_{*}^{2}}}{p_{E}^{2}+m^{2}},$ (35)
where the overall factor of $i$ has been absorbed into the continuation of the Green’s function to Euclidean signature. Comparing (30) with (35), we see that the factor $e^{-p^{2}/M_{*}^{2}}$ in Minkowski space has become $e^{-p_{E}^{2}/M_{*}^{2}}$ in Euclidean space. The crucial point is that $p_{E}^{2}$ is positive and grows without bound as $|p_{E}^{\mu}|\to\infty$, so the exponential factor provides strong UV damping:
$e^{-p_{E}^{2}/M_{*}^{2}}\sim\begin{cases}1,&p_{E}^{2}\ll M_{*}^{2},\\ \exp(-p_{E}^{2}/M_{*}^{2}),&p_{E}^{2}\gg M_{*}^{2}.\end{cases}$ (36)
This ensures that (35) converges absolutely in the UV, even in theories that would be power-counting nonrenormalizable without the regulator.
The argument generalizes straightforwardly to higher-loop diagrams and to other choices of the entire function $F(z)$. In all cases, the momentum dependence of the regulator in Euclidean space is obtained by the replacement $z=\Box/M_{*}^{2}\to-p_{E}^{2}/M_{*}^{2}$ in the eigenvalue relation. For example, if $F(z)=e^{-z}$, then $F(\Box/M_{*}^{2})$ would contribute a factor $e^{+p_{E}^{2}/M_{*}^{2}}$ and would spoil UV convergence; such choices are therefore excluded. The physically relevant class of regulators is characterized by entire functions $F(z)$ that decay sufficiently fast along the positive real axis $z=p_{E}^{2}/M_{*}^{2}$.
IV Gauge invariance and the background-field method
So far we have considered only a scalar field with no gauge interactions. In HUFT and in more general gauge theories, we require the regulator to respect gauge invariance and BRST symmetry. This is achieved by constructing the regulator as an entire function of the covariant operator $\Box=-g^{\mu\nu}D_{\mu}D_{\nu}$, rather than of the ordinary Laplacian.
In the background-field formalism, one splits the gauge field $A_{\mu}$ into a classical background part $A_{\mu}^{\text{bg}}$ and a quantum fluctuation $a_{\mu}$:
$A_{\mu}=A_{\mu}^{\text{bg}}+a_{\mu}.$ (37)
One chooses a background-covariant gauge-fixing condition for $a_{\mu}$ that preserves invariance under gauge transformations of the background field $A_{\mu}^{\text{bg}}$. The effective action $\Gamma[A_{\mu}^{\text{bg}}]$ obtained after integrating out $a_{\mu}$ is then manifestly gauge invariant under transformations of the background. In this setting, the regulator takes the form:
$F\!\left(\frac{\Box_{\text{adj}}}{M_{*}^{2}}\right),$ (38)
where $\Box_{\text{adj}}$ is the covariant Laplace–Beltrami operator acting in the adjoint representation on the quantum gauge fluctuation $a_{\mu}$ and on the ghost fields. Because $\Box_{\text{adj}}$ transforms covariantly under background gauge transformations, the regulated kinetic terms and interaction vertices inherit the same symmetry. The resulting Feynman rules involve propagators and vertices multiplied by form factors $F(-p^{2}/M_{*}^{2})$ in Minkowski momentum space, as in the scalar case.
Expanding around the perturbative vacuum ($g_{\mu\nu}\to\eta_{\mu\nu}$, $A_{\mu}^{\text{bg}}\to 0$), $\Box_{\text{adj}}$ reduces to $\Box_{M}$ acting componentwise on the gauge and ghost fields. The plane-wave eigenfunctions and eigenvalues are therefore identical to the scalar case, and the transition to Euclidean momentum space proceeds exactly as in Section 4. The Ward and Slavnov–Taylor identities are preserved because the regulator is constructed from background-covariant operators, and the exponential damping factors $e^{-p_{E}^{2}/M_{*}^{2}}$ appear only as multiplicative factors that do not introduce additional poles or branch cuts.
V Liouville’s theorem and the singularity at infinity
We address the concern raised by Liouville’s theorem. Liouville’s theorem states that any bounded entire function must be constant, so any nonconstant entire function must be unbounded and has an essential singularity at the point $z=\infty$ on the Riemann sphere [27, 28].
In our context, $z$ is effectively identified with the complexified eigenvalue of $\Box/M_{*}^{2}$, which in the perturbative vacuum is $p^{2}/M_{*}^{2}$ in Minkowski space or $p_{E}^{2}/M_{*}^{2}$ in Euclidean space. The fact that $F$ is unbounded as $|z|\to\infty$ simply reflects the fact that $F$ must grow in certain directions in the complex plane. However, the physical amplitudes are constructed from the values of $F(z)$ along specific directions such as the real Euclidean axis $z=-p_{E}^{2}/M_{*}^{2}\leq 0$, where $p_{E}^{2}$ is real and nonnegative, the real Minkowski $p^{2}$-axis, approached with the standard $i\epsilon$-prescription for timelike momenta. By choosing $F(z)$ to be free of zeros in the finite complex plane and to have suitable decay along the Euclidean axis, we guarantee that the propagators do not acquire additional poles or branch cuts in the finite $p^{2}$-plane; the only singularities correspond to the physical particle masses. Loop integrals are exponentially damped along the Euclidean axis, yielding UV-finite amplitudes.
The essential singularity at $z=\infty$ is never encountered by any physically relevant contour in the complex $p^{2}$-plane. In particular, the analytic continuation of Green’s functions is controlled by the standard Feynman prescription and by the location of branch cuts associated with multi-particle thresholds, none of which are altered by the entire-function regulator.
Thus, Liouville’s theorem constrains the global behaviour of $F(z)$ on the entire complex plane, but does not obstruct the existence of well-defined, UV-finite, and unitary quantum field theories (QFTs) regulated by such entire functions. The key property we exploit is the absence of zeros and poles of $F(z)$ at finite $z$, not the behaviour at infinity. The essential singularity exists at infinity, but it exists in parts of the complex plane that are not probed in any physical processes.
VI Nonlocality, microcausality, and the local limit
The operator $F(\Box/M_{*}^{2})$ is nonlocal in position space, since it involves infinitely many derivatives. It is instructive to rewrite it as a convolution with a kernel K(xy)K(x-y). In flat space, acting on a scalar field, we may write:
$F\!\left(\frac{\Box_{M}}{M_{*}^{2}}\right)\phi(x)=\int d^{4}y\,K(x-y)\,\phi(y),$ (39)
where the kernel K(xy)K(x-y) is obtained by Fourier transforming the momentum-space form factor:
$K(x-y)=\int\frac{d^{4}p}{(2\pi)^{4}}\,F\!\left(-\frac{p^{2}}{M_{*}^{2}}\right)e^{-ip\cdot(x-y)}.$ (40)
For the choice $F(z)=e^{z}$, we have:
$K(x-y)=\int\frac{d^{4}p}{(2\pi)^{4}}\,e^{-p^{2}/M_{*}^{2}}e^{-ip\cdot(x-y)}.$ (41)
Upon Wick rotation to Euclidean space, this becomes the Fourier transform of a Gaussian:
$K_{E}(x_{E}-y_{E})=\int\frac{d^{4}p_{E}}{(2\pi)^{4}}\,e^{-p_{E}^{2}/M_{*}^{2}}e^{ip_{E}\cdot(x_{E}-y_{E})},$ (42)
which is itself a Gaussian in $(x_{E}-y_{E})^{\mu}$ with characteristic width of order $M_{*}^{-1}$. The nonlocality is controlled by the length scale:
$\ell_{*}\sim M_{*}^{-1}.$ (43)
Fields at $x$ are smeared over a neighbourhood of size $\ell_{*}$ in spacetime.
Microcausality in QFT is usually formulated in terms of commutators of local operators at spacelike separation. In the presence of nonlocal regulators of this type, equal-time commutators develop exponentially suppressed tails outside the light cone, controlled by the scale $\ell_{*}$. In the limit $M_{*}\to\infty$, the kernel K(xy)K(x-y) becomes sharply peaked at $x=y$ and tends to a Dirac delta function, so that
$F\!\left(\frac{\Box}{M_{*}^{2}}\right)\phi(x)\xrightarrow[M_{*}\to\infty]{}\phi(x),$ (44)
and locality is recovered. For finite but large $M_{*}$, the theory is quasi-local, with deviations from strict microcausality confined to scales $\lesssim\ell_{*}$. In HUFT, $M_{*}$ is taken to be near the Planck or GUT scale, so these effects are negligible at accessible energies and distances.
Importantly, the nonlocality introduced by $F(\Box/M_{*}^{2})$ does not alter the particle content of the theory. The pole structure of the propagators remains unchanged, and the exponential factors only suppress high-momentum modes. As a result, the spectral representation of two-point functions and the optical theorem are preserved, ensuring unitarity.
VII Paley–Wiener bounds and quasi-locality
The relation between entire-function regulators in momentum space and quasi-locality in position space can be formulated more precisely in terms of Paley–Wiener–type theorems [29, 30, 31]. In flat space, acting on a scalar field, we have:
$F\!\left(\frac{\Box_{M}}{M_{\ast}^{2}}\right)\,\phi(x)=\int d^{4}y\,K(x-y)\,\phi(y),$ (45)
where the kernel is the Fourier transform of the momentum-space form factor:
$K(x-y)=\int\frac{d^{4}p}{(2\pi)^{4}}\,F\!\left(-\frac{p^{2}}{M_{\ast}^{2}}\right)e^{-ip\cdot(x-y)}.$ (46)
For the choice $F(z)=e^{z}$, we find explicitly that, after Wick rotation to Euclidean space, K(xy)K(x-y) is a Gaussian with characteristic width $\ell_{\ast}\sim M_{\ast}^{-1}$, so that fields at $x$ are smeared over a neighbourhood of size $\ell_{\ast}$ in spacetime and locality is recovered in the limit $M_{\ast}\to\infty$.
Paley–Wiener theorems provide the corresponding “if and only if” statement, that an entire function $F$ of the Laplace operator whose growth in complex momentum space is of finite exponential type is equivalent to a position-space kernel K(xy)K(x-y), which is not compactly supported but decays at most exponentially fast at large spacelike separation. So, once one demands an entire, UV-damping form factor $F(-p^{2}/M_{\ast}^{2})$ with good behaviour along the Euclidean axis, the presence of exponentially small but nonzero tails in K(xy)K(x-y) is Inevitable; strict compact support of the kernel would be incompatible with nontrivial entire UV suppression.
In particular, the quasi-local behaviour discussed above in which the commutators of smeared operators at spacelike separation are suppressed as $\exp(-\alpha M_{\ast}\rho)$ up to polynomial factors in the geodesic distance $\rho$ is precisely of Paley–Wiener type. The scale $M_{\ast}^{-1}$ controls the onset of nonlocality in momentum space and also controls the decay length of the nonlocal kernel in position space. The local limit $M_{\ast}\to\infty$ corresponds to sending the exponential type of $F$ to zero.
Given HUFT’s structural assumptions of an entire function regulator with a proper-time representation and the Paley–Wiener–type locality bounds, the admissible regulators form a Paley–Wiener class of entire form factors that exhibit Gaussian-type exponential damping at large Euclidean momentum, up to subleading exponentially suppressed corrections and redefinitions of the nonlocality scale $M_{\ast}$. Our choice $F(z)=e^{z}$ is the simplest representative of this class. From the Paley–Wiener theorem, we learn that the regulator must lie in the class of Gaussian entire functions with asymptotic behaviour of $e^{-}p^{2}_{E}/M_{*}^{2}$, up to subleading polynomial factors and redefinitions of M*
VIII Discussion and outlook
We have given a detailed and fully covariant derivation of how gauge-invariant entire-function regulators operate in HUFT and related nonlocal QFTs. The essential points can be summarized as follows: the regulator is implemented as an entire function $F(\Box/M_{*}^{2})$ of the covariant Laplace–Beltrami operator $\Box=-g^{\mu\nu}D_{\mu}D_{\nu}$, where $D_{\mu}$ is the gauge- and diffeomorphism-covariant derivative and $M_{*}$ is the nonlocality scale. In the perturbative vacuum, where $g_{\mu\nu}\to\eta_{\mu\nu}$ and $A_{\mu}\to 0$, the operator $\Box$ reduces to the Minkowski d’Alembertian $\Box_{M}$, which is diagonalized by plane waves with eigenvalues $p^{2}$. Acting on a plane wave, the entire operator $F(\Box_{M}/M_{*}^{2})$ reduces to multiplication by $F(-p^{2}/M_{*}^{2})$. After Wick rotation, this becomes $F(-p_{E}^{2}/M_{*}^{2})$ in Euclidean momentum space, providing exponential UV damping of loop integrals for appropriate choices of $F$. Gauge invariance and BRST symmetry are preserved by constructing the regulator from background-covariant operators in the background-field formalism. The Ward and Slavnov–Taylor identities retain their standard form, and no unphysical degrees of freedom are introduced. Liouville’s theorem implies that nonconstant entire functions are unbounded and possess an essential singularity at infinity, but this has no physical consequences. Only the finite complex $p^{2}$-plane and the real Euclidean axis are probed by perturbative amplitudes, where $F(-p_{E}^{2}/M_{*}^{2})$ is analytic and free of zeros. The nonlocality associated with $F(\Box/M_{*}^{2})$ corresponds to a smearing of fields over a length scale $\ell_{*}\sim M_{*}^{-1}$. Locality and microcausality are recovered in the limit $M_{*}\to\infty$, and for large but finite $M_{*}$ the theory is quasi-local with negligible deviations at accessible scales.
Moreover, the same gauge- and diffeomorphism-covariant entire-function regulator $F(\Box/M_{*}^{2})$ can be applied to the gravitational sector by dressing curvature invariants with $F(\Box/M_{*}^{2})$. In Euclidean momentum space this yields an exponentially suppressed graviton propagator $\Delta_{\rm grav}(p_{E}^{2})\sim e^{-p_{E}^{2}/M_{*}^{2}}/p_{E}^{2}$, so that graviton loop integrals become UV convergent without introducing additional ghost poles. Thus, the mechanism that makes the gauge and matter sectors finite naturally extends to quantum gravity, rendering the full framework renormalizable and plausibly finite, to all orders.
In the specific context of HUFT, these results provide a rigorous foundation for the use of entire-function regulators to render both gauge and gravitational sectors finite while maintaining a well-defined particle spectrum and preserving the geometric structure of the theory. They also clarify why Liouville’s theorem, often cited as an objection at a heuristic level, does not pose an obstacle to the physical viability of such regulators.